OpenStudy (anonymous):

What are the exact solutions of x2 − 3x − 7 = 0?

OpenStudy (anonymous):

use the quadratic frmula x =[ --3 +- sqrt(9 - 4 * 1*-7)] / 2 =[ 3 + sqrt37] /2 or [3 - sqrt37] / 2

jimthompson5910 (jim_thompson5910):

\[\Large x=\frac{-B\pm\sqrt{B^2-4AC}}{2A}\] \[\Large x=\frac{-\left(-3\right)\pm\sqrt{\left(-3\right)^2-4\left(1\right)\left(-7\right)}}{2\left(1\right)}\] \[\Large x=\frac{3\pm\sqrt{\left(-3\right)^2-4\left(1\right)\left(-7\right)}}{2\left(1\right)}\] \[\Large x=\frac{3\pm\sqrt{9-4\left(1\right)\left(-7\right)}}{2\left(1\right)}\] \[\Large x=\frac{3\pm\sqrt{9-\left(-28\right)}}{2\left(1\right)}\] \[\Large x=\frac{3\pm\sqrt{9+28}}{2\left(1\right)}\] \[\Large x=\frac{3\pm\sqrt{37}}{2\left(1\right)}\] \[\Large x=\frac{3\pm\sqrt{37}}{2}\] \[\Large x=\frac{3+\sqrt{37}}{2} \ \text{or} \ x=\frac{3-\sqrt{37}}{2}\]